OpenStudy (flygirlmandi):

evaluate -5g-6 for g=-2 a-4 b-16 c-10 d- -13 thanks :)

OpenStudy (alexandervonhumboldt2):

-5g-6 substitute: -5*(-2)-6 -*-=+ -5*(-2)-6=5*2-6=10-6=??? use PEDMAS

OpenStudy (flygirlmandi):

@AlexandervonHumboldt2 is it 4 ....its 4 right ?

OpenStudy (alexandervonhumboldt2):

it is 4

OpenStudy (alexandervonhumboldt2):

wait

OpenStudy (alexandervonhumboldt2):

what idiot made this question?

OpenStudy (alexandervonhumboldt2):

there are no corret choices

OpenStudy (flygirlmandi):

it was for my school ?

OpenStudy (alexandervonhumboldt2):

there must be a typo

OpenStudy (phi):

I'm thinking those are dashes, not minus signs. so choice a

OpenStudy (alexandervonhumboldt2):

maybe :)

OpenStudy (flygirlmandi):

im going with 4....thanks guys :)

OpenStudy (alexandervonhumboldt2):

np :)